OpenStudy (anonymous):

which of the following is the binomial factor of the factorization of m15-1,000?

OpenStudy (anonymous):

What are the choices?

OpenStudy (anonymous):

m^5-10 m^5+10 m^5+100 m^3+10

OpenStudy (anonymous):

That's m^15 - 1000 right? Then the answer would be the first one I believe. I'm not positive, but I think it's right.

OpenStudy (anonymous):

okay thank you very much

OpenStudy (anonymous):

you were right

OpenStudy (anonymous):

it is the first one as you have the difference of two cubes \[a^3-b^3=(a-b)(a^2+ab+b^2)\] with \(a=m^3,b=10\)